OpenStudy (anonymous):

Find 3 consecutive integers such that the product of the two smaller integers is 38 more than twice the largest. :)

OpenStudy (anonymous):

If the integers are consecutive, then you only need one variable to describe each of them: \(x, x+1,x+2\). The product of the two smaller ones is 38 more than twice the largest. The product of the two smaller ones is \[(x)(x+1)\] and 38 more than twice the largest is \[2(x+2)+38\]

OpenStudy (anonymous):

z-y = 1 y-x = 1 xy = 2z + 38 y= x+ 1 z = 2 + x sub these into third equation x(x+1) = 2(2+x) + 38 x^2 + x = 4 + 2x + 38 x^2 - x -42 = 0 x = 6 so three numbers are 6, 7 and 8

OpenStudy (anonymous):

yep, I completely misread the problem >.<